Produce graphs of f that reveal all the important aspects of the curve. Then use calculus to find the following. (Enter your answers using interval notation. Round your answers to three decimal places.)

f(x) = ex − 0.196x4

Find the inflection points of the function.

(x, y) =

(smallest value)

(x, y) =
(x, y) =     (largest x-value)



Find the interval(s) where the function is concave up.

Find the interval(s) where the function is concave down.
Answer #1

e 0.19624 Given f(x) = at f00)=0 Inflection points occors f(x) = d da ex - 0.196 (423) -e da x=n-an- - ex 0.784 03 FCO) d dx = 1.171 - 2.8586 2.859 = (1.17) inflection points are (-0.506, 0.590 ) /smallest se value I female (1.171, 2.859 ) loogest